Frm a uniform circular disk of mass M and radius R a small cirucular disk of radius R/2is removed in such a way that both hav a common tangent.find the distance of centre of mass of remaining part frm the centre of the disk.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From unitary method fidn the mass of the cut part that would be = M/4.
And assum -ve mass be added to the remaining part that is -M//4 to complete the disc.
Therefore,
Ycom = (M*0 + (-M/4)*R/2)/(3M/4) = -R/6
and
XCOM =0
R/6 is the distance from centre of the disk.